I been working on this problem all night and I am stuck. The book says the answer is 283.4 m, but I dont know how to get it. Thanks for your help
A man runs 300 m north and then 400 m to the southeast (at a 45 degree angle). If he has run at a constant speed, how far away is he from the starting position?

Answer by edjones(8007)

s-a-s triangle
a^2=b^2+c^2-2bc*cos(a)
=300^2+400^2-2*300*400*cos(45)
=90000+160000-120000sqrt(2)
=80294.36
a=283.4 m
.
Ed